when four times a certain number is increased by three, the result is twenty-one more than the number. Find the number.
Answer by mananth(16949)

when four times a certain number is increased by three, the result is twenty-one more than the number. Find the number.
Let the number be x
4 times the number = 4x
increased by 3 = 4x+3
221 more than the number = x+21
4x+3=x+21
4x-x=21-3
3x=18
x=6